The following transactions and adjusting entries were completed by a paper-packaging company called Gravure Graphics International during 2018 and 2019. The company uses straight-line depreciation for trucks and other vehicles, double-declining-balance depreciation for buildings, and straight-line amortization for patents.

2018
January 2 Paid $88,000 cash to purchase storage shed components.
January 3 Paid $4,000 cash to have the storage shed erected. The storage shed has an estimated life of 10 years and a residual value of $7,000.
April 1 Paid $31,000 cash to purchase a pickup truck for use in the business. The truck has an estimated useful life of five years and a residual value of $5,000.
May 13 Paid $300 cash for minor repairs to the pickup truck's upholstery.
July 1 Paid $27,000 cash to purchase patent rights on a new paper bag manufacturing process. The patent is estimated to have a remaining useful life of five years.
December 31 Recorded depreciation and amortization on the pickup truck, storage shed, and patent.
2019
June 30 Sold the pickup truck for $26,000 cash. (Record the depreciation on the truck prior to recording its disposal.)
December 31 Recorded depreciation on the storage shed. Also determined that the patent was impaired and wrote off its remaining book value (i.e., wrote down the book value to zero).

Required:

Prepare the journal entries required on each of the above dates. (If no entry is required for a transaction/event, select "No Journal Entry Required" in the first account field. Do not round intermediate calculations.)
Answer #1

Answer :

Date Account title Debit credit
2018
Jan 2 storage shed components 88000
cash 88000  
Jan 3 storage shed components 4000
cash 4000
April 1 Truck 31000
cash 31000
may 13 Repair and maintenance expense 300
cash 300
July 1 Patent 27000
cash 27000
December 31 Depreciation expense -storage shed components 8500
Depreciation expense -Truck 3900
Accumulated depreciation-storage shed components 8500
Depreciation expense -Truck 3900
December 31 Amortisation expense-patent 2700
Accumulated amortisation 2700
2019
June 30 Depreciation expense -truck 2600
Accumulated depreciation -truck [5200*6/12] [jan-june] 2600
June 30 cash 26000
Accumulated depreciation [3900+2600]] 6500
Gain on sale of truck 1500
Truck 31000
Dec 31 Depreciation expense 8500
Accumulated depreciation -storage shed components 8500
Dec 31 Impairment loss 24300
Accumulated amortisation 2700
Patent 27000

Depreciation for 2018 :storage shed = [cost -salvage value ]/useful life

                      =[(88000+4000)-7000]/10

                      = [92000 -7000]/10

                     = 85000/10

                       =8500

Depreciation expense for 2018 :Truck = [31000-5000]/5

                       = 5200

depreciation for part months of 2018 (april - dec ]= 5200*9/12= 3900

Amortisation expense - patent =[27000/5]=5400*6/12 = 2700